OpenStudy (anonymous):

what is the derivative of sin^2x

OpenStudy (yttrium):

Do you know how to apply the power rule? :)

OpenStudy (yttrium):

It's like \[\frac{ d }{ dx } \sin^n (x) = nsin ^{n-1}x (dx)\]

OpenStudy (anonymous):

@josh7 its=2sinx

OpenStudy (anonymous):

\[(\sin ^{2}x)^{'} =\] 2sinxcosx

OpenStudy (anonymous):

u=sinx (sin^2x)'=f(u)'u'